OpenStudy (azureflame241):

After working to Simplify x^(3/2)*(x+x^(5/2)-x^2), i got x^(11/2)... can someone check to make sure if this is correct?

OpenStudy (anonymous):

those fraction exponents will be change in radical . like this \[\sqrt[2]{x ^{3}}\]

zepdrix (zepdrix):

They're much easier to work with as fractions actually. :o\[\Large x^{3/2}(x+x^{5/2}-x^2)\]Distrubuting the x^3/2 gives us,\[\large x^{(1+3/2)}+x^{(5/2+3/2)}-x^{(2-3/2)}\]badaboom,\[\large x^{5/2}+x^{8/2}-x^{1/2}\]

zepdrix (zepdrix):

Ah that last term should be x^(2+3/2), my bad!\[\large x^{(1+3/2)}+x^{(5/2+3/2)}-x^{(2+3/2)}\]

zepdrix (zepdrix):

So I guess that last term becomes,\[\Large x^{7/2}\]

OpenStudy (azureflame241):

ya thats what i got when i was working on it

zepdrix (zepdrix):

Hmm how did you end up with x^11/2? D: make a boo boo somewhere?

OpenStudy (azureflame241):

ya its a mistake...it should be x^(6/2) ===> x^3

zepdrix (zepdrix):

Careful! You can't add the terms together like that. We do NOT have, \[\Large \frac{5}{2}x+\frac{8}{2}x-\frac{7}{2}x\]Which we would be able to add. We have different powers of x, which cannot be combined through addition.\[\Large x^{5/2}+x^{8/2}-x^{7/2}\]That would be our final answer.

zepdrix (zepdrix):

@azureflame241 You could simplify the middle term to x^4 though if you wanted to. :)